Shift summary for evening shift, 06/06/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Wed, 9 Jun 1999 22:21:42 -0700 (MST)

Shift Summary for evening shift, 06/06/99

On Shift: Ashkan, Sydney

Also: Ed, RickK, Hogan, Miguel, Patty, Elizabeth

Hours of beam delivered at all intensities: 2
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1

Accounting of lost time:
TeV problems, NM2 beam studies, blown fuse in crate 27 (DC4Y).

========================================================================
Runs ended this shift:
Run 13466 COSMICS
Run 13472 - 200K events - BEAM_MUON_4E12

Tapes brought to FCC this shift: None.

UNVAULTED
QKC001-3
QKS001-3
QKK001-3
QKB001
after Arun discovers that the magnetic labels are incorrect.
An investigation is underway.

The current tapes
QKC004-6
QKS004-6
QKK004-6
QKB002
were checked and found to be correctly labelled.

========================================================================
Accesses during shift:
Ashkan, Miguel, and Hogan replace blown fuse in crate 27 (DC4Y).

========================================================================
Overview of shift:
RickC and RickK finish target scan. DC crew on hand
to do DC plateaus and studies with muons but no beam for most of shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13466 15:30 COSMICS run in progress.
No beam: large LCW leak. No estimate.

16:00 Elizabeth changes some REG HVs

16:50 Beam is back. RickF continues beam
studies with mesons.

17:00 RickK reduces SA & CIA thresholds to
300mV. This will take effect after run 13466.

18:10 Stop run 13466.

19:00 RickK changes B02 prescale from 7 to 9
to reduce Ke3 output and improve tape
volume balancing.

20:10 To improve tape volume balancing, Sydney changes
the BEAM_MUON configuration to send MUON runs to
QKS (instead of QKK) tapes.

20:55 RickF is done with his beam studies.
No beam. No estimate.

13472 21:30 Beam is back. Ask MCR to set NMS2S2 = 500,
NMS2S3 = 100 for DC muon studies. Start
MUON run 13472.

13473 22:00 Stop run 13472. Start MUON run 13473.

22:40 Ashkan, Miguel, and Hogan go on access to
replace blown fuse in CAMAC crate 27 (DC4Y).

22:45 CH13: TeV quench at B4.

23:55 Beam is back.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

After plateauing DCs, DC crew will take ~2-3 million muons. We
will then switch to BEAM_832 running at 2E12. The following
procedure should be followed in switching from muons to hadrons:

o Bring up the analysis magnet at the current polarity
o Call MCR to change NM2S2/NM2S3 back to 1500/875.
o After verifying with epicure that NM2S2/NM2S3 are at the above settings..
o Pull out exp. crash switch
o Call MCR to open beam stop